40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 01:40:37 2013, Manasa, Update, Alignment, Final IFO alignment- in progress 
    Reply  Wed Feb 20 03:13:10 2013, yuta, Update, Alignment, clipping centering checklist ClippingCenteringChecklist.pdf
       Reply  Wed Feb 20 20:58:37 2013, yuta, Update, Alignment, clipping centering checklist ClippingCenteringChecklist.pdf
          Reply  Wed Feb 20 21:12:37 2013, rana, Update, Alignment, clipping centering checklist 
             Reply  Wed Feb 20 21:56:08 2013, yuta, Update, Alignment, clipping centering checklist 
    Reply  Wed Feb 20 10:13:41 2013, yuta, Update, Alignment, PRC flashing brighter than last week 
Message ID: 8113     Entry time: Wed Feb 20 01:40:37 2013     Reply to this: 8114   8115
Author: Manasa 
Type: Update 
Category: Alignment 
Subject: Final IFO alignment- in progress 

[Yuta, Sendhil, Jamie, Jenne, Rana]

1. After the MC centering, we tried to align the IFO using IPPOS and IPANG as reference. This did not recover the alignment perfectly. We were clipping at the BS aperture. Using TTs, we centered the beam at BS and PRM.
2. Using TTs, the beam was centered at ITMY and ETMY.
3. IPPOS and IPANG mirrors in-vacuum were aligned and were centered at the out-of-vacuum optics.
4. We checked the centering of the beam on optics in the BS and ITMY chamber. (Yuta will make an elog with the layout)
5. We retro-reflected ITMY at the BS and aligned ETMY such that we saw a couple of bounces in the arm cavity.
6. Using BS, the beam was steered to go through the center of ITMX and ETMX.
7. At this point we were able to see the MI fringes at the AS port.
8. We made fine alignments to the ITMX such that we saw MI reflected at the Farday.
9. We retro-reflected ITMX and aligned ETMX to see the beam bounce at the ITMX.
10. We aligned PRM such that PRC flashes. But we were not happy with the flashes (they were in higher order modes). We suspect that minor tuning of the input pointing might be necessary.
11. We closed for the day

ELOG V3.1.3-